OpenStudy (mallorysipp234):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)? g(x) − f(x) = 5x − 8 − 3x g(x) − f(x) = 5x − 12 − 3x g(x) − f(x) = 3x − 8 − 5x g(x) − f(x) = 3x + 12 − 5x

OpenStudy (cggurumanjunath):

3x+2-5x+10 =g-f

OpenStudy (cggurumanjunath):

is it clear @mallorysipp234 ?

OpenStudy (mallorysipp234):

so it would be D? @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

:)

OpenStudy (cggurumanjunath):

yes.

OpenStudy (mallorysipp234):

can you help me with a another?

OpenStudy (mallorysipp234):

If f(x) = log2 (x + 4), what is f ^ −1(3)? 0 2 4 8